# Efficient Solutions for Mitigating Initialization Bias in Unsupervised Self-Adaptive Auditory Attention Decoding

### v1

@@ 3.1 Dataset and Hyperparameters
We evaluate all methods on a public dataset from [13] used in [7, 8]. It contains 72-min EEG recordings from 16 normal-hearing subjects attending to one of two competing speakers at $\pm 90^{\circ}$ azimuth, with corresponding audio data. Following [7], audio signals are processed using a gammatone filterbank, envelopes are extracted via power-law operation (exponent 0.6) and summed across subbands. Both EEG and speech envelopes are filtered to 1-9 Hz, downsampled to 20 Hz, and cut into 60-s segments.
@@ 3.2 Results
We evaluate decoding performance in both transductive and inductive settings. In the transductive setting, predictions are generated for the data on which the unsupervised model is trained, while inductive decoding assesses model generalization to unseen data. We also report normalized CPU time, defined as the ratio of each method’s computational time to that of the baseline single-encoder approach from Section 2.1. To examine the performance under limited training data settings, random 3-fold cross-validation is used, with the training sets subsampled to target durations. Train/test splits are identical across methods.
@@ 4 Conclusion
We proposed three computationally efficient solutions to mitigate initialization bias in unsupervised self-adaptive AAD. The two-encoder version trains encoders for both attended and unattended features. The soft version replaces hard segment assignments with probabilistic weights. The sum-initialized single-encoder method initializes the model with a composite signal. For smaller datasets, the sum-initialized approach is the top performer while matching the baseline’s computational cost. With larger datasets, the soft-label method becomes competitive, approaching the cross-validated variant’s accuracy at low cost. A limitation of this work is that evaluation was restricted to a single dataset. Future validation on more diverse datasets is needed to confirm generalizability.

### v2

@@ 3.1 Dataset and Hyperparameters
We evaluate [^1] For conciseness, we report results for a single dataset. Similar results for an additional dataset are available in the repository linked in the abstract. all methods on a public dataset from [2] used in [6, 10]. It contains 72-min 64-channel EEG recordings (all channels used) from 16 normal-hearing subjects attending to one of two competing speakers at $\pm 90^{\circ}$ azimuth, with corresponding audio data. Following [6], audio signals are processed using a gammatone filterbank, envelopes are extracted via power-law operation (exponent 0.6) and summed across subbands. Both EEG and speech envelopes are filtered to 1-9 Hz [9], downsampled to 20 Hz, and cut into 60-s segments.
@@ 3.2 Results
We evaluate decoding performance in both transductive and inductive settings. In the transductive setting, predictions are generated for the data on which the unsupervised model is trained, while inductive decoding assesses model generalization to unseen data. We also report normalized CPU time, defined as the ratio of each method’s computational time (Windows 11, Intel Core i7-13700F, single thread) to that of the baseline single-encoder approach from Section 2.1. To examine the performance under limited training data settings, random 3-fold cross-validation is used, with the training sets subsampled to target durations. Train/test splits are identical across methods.
@@ 4 Conclusion
We proposed three computationally efficient solutions to mitigate initialization bias in unsupervised self-adaptive AAD. The two-encoder version trains encoders for both attended and unattended features. The soft version replaces hard segment assignments with probabilistic weights. The sum-initialized single-encoder method initializes the model with a composite signal. For smaller datasets, the sum-initialized approach is the top performer while matching the baseline’s computational cost. With larger datasets, the soft-label method becomes competitive, approaching the cross-validated variant’s accuracy at low cost.